OpenStudy (anonymous):

could someone please help me solve this problem? find the exact value, in radians, without using calculator, of sin -1(square of 3/-2)

OpenStudy (amistre64):

how to solve? its more memorization than anything; these numbers are most common with the 60-30 degrees

OpenStudy (amistre64):

sin(t) = sqrt(3)/-2

OpenStudy (amistre64):

-60 = -pi/3

OpenStudy (mathteacher1729):

http://upload.wikimedia.org/wikipedia/commons/4/4c/Unit_circle_angles_color.svg that is your best friend. "What value of theta give sin(theta) = -sqrt(3)/2" is what your question is asking. :)

OpenStudy (amistre64):

sin-1 has a range of -pi/2 to pi/2; so -pi/3 is the value it gives